Let's assume that your baumol-tobin model is $ 54,000 a year. The annual interest rate of the savings deposit account is 2.5% and we assume that you keep your income in your account.

1- Assume you go to the bank 4 times a month, how many times do you go to the bank in total?

2- How much money do you keep on average for a year?

3- Now, every time you go to the bank, let F cost $ 0.42. how much of your income do you have??
Answer #1

1. The total no of times to the bank is 12×4 =48. 12 is no of months in a year.

2. Money to be kept on an average in a year is:

$54000/(2*48) = $562.5

3.The total (opportunity) cost equals

Total Cost = Cost from foregone interest + Cost of trips to the bank = i(y/2n) + F*N = 0.025*($54000/96) + 0.42*48 = $34.2225